HI, I need help with answering these questions. Please explain and answer all parts. Data for all the countries and then the question at the bottom.

Sugar Consumption Per Capita.csv Country Albania Argentina Armenia Australia Austria Azerbaijan Belarus Belgium Bosnia and Herzegovina 13.4 4754.197861 Europe Brazil Canada Chile China Colombia Czech Republic Denmark Egypt Estonia Finland France Georgia Germany Ghana Greece Hungary Iceland India Indonesia Iran Sugar GDP Continent 15.3 4556.144342 Europe 38.1 13693.70379 South America 33.2 3421.704509 Europe 34.1 62080.98242 Europe 37.949485.48219 Europe 13.9 7189.691229 Europe 31.8 6305.773662 Europe 41.4 46463.60378 Europe 36.5 12576.19559 South America 31.3 51790.56695 North America 41.7 14510.9661 South America 6.2 5447.309378 Asia 23.2 7124.54892 South America 30.6 20584.92655 Europe 38 59911.90466 Europe 26.4 2972.583516 Africa 31.4 16982.30031 Europe 25 48694.53514 Europe 31.9 42578.17709 Europe 27 3219.605869 Europe 34 44354.68494 Europe 9.6 1594.030809 Africa 22.6 26061.44027 Europe 13.1 13784.18353 Europe 33.9 44019.3907 Europe 18.3 1539.606447 Asia 12.5 3469.753726 Asia 26.1 7006.047183 Asia
Ireland Israel Italy Japan Kazakhstan Kenya Kyrgyzstan Latvia Lithuania Luxembourg Macedonia Malaysia Malta Mexico Moldova Mongolia Montenegro Morocco Netherlands New Zealand Nigeria Norway Pakistan Peru Philippines Poland Portugal Romania Russia Serbia Slovenia South Africa 26.3 49387.27333 Europe 29.133250.5065 Asia 25 36988.16405 Europe 16 46203.69804 Asia 25.8 11357.94549 Asia 14.3 816.441535 Africa 20.2 1123.883168 Asia 38.8 13827.36026 Europe 28.7 14227.68554 Europe 10.8 111913.1844 Europe 30.7 4940.953345 Europe 38.4 10058.04302 Asia 44.1 22346.31841Europe 33.3 9802.894353 South America 24 1970.840003 Europe 11.3 3181.104401 Asia 18.3 7253.359249 Europe 36 3044.107888 Africa 40.649886.28451 Europe 44.3 37192.64543 Australia 8.6 2507.682969 Africa 28 99091.0945 Europe 23.1 1212.978046 Asia 19.3 5759.398972 South America 19.4 2357.570924 Asia 40.6 13384.78217Europe 21.7 22532.50772 Europe 18.9 9063.676031 Europe 42.4 13324.28784 Europe 25.7 6047.73788 Europe 17 24478.31997 Europe 30.3 7830.505168 Africa
South Korea Spain Sweden Switzerland Tajikistan Thailand Turkey Turkmenistan Ukraine United Kingdom USA Uzbekistan Venezuela Vietnam 17.4 24155.8293 Asia 21.5 31117.89747 Europe 33.7 56724.36224 Europe 49.2 83270.24474 Europe 16.4 834.6586566 Asia 35.5 5192.118907Asia 27.3 10604.5524Asia 7.3 5724.541586 Asia 39.6 3575.490646 Europe 35.8 38927.06931 Europe 28.1 49854.52267 North America 9.3 1544.827773 Asia 34.8 10727.98209 South America 9.6 1543.02695 Asia
Question 8: Calculate and interpret the correlation between sugar consumption and GDP. Describe your findings in one paragraph or less.
Answer #1

Sugar Values X is
∑ = 1994.9
Mean = 26.599
∑(X - Mx)2 = SSx = 8258.01

GDP Values Y is
∑ = 1626334.401
Mean = 21684.459
∑(Y - My)2 = SSy = 42290592466.331

X and Y Combined
N = 75
∑(X - Mx)(Y - My) = 5615579.317

R Calculation
Correlation r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

Correlation r = 5615579.317 / √((8258.01)(42290592466.331)) = 0.3005

Meta Numerics (cross-check)
Correlation r = 0.3005

Although technically a positive correlation, the relationship between your variables is weak (nb. the nearer the value is to zero, the weaker the relationship).

Click here to calculate a p-value.

The value of R2, the coefficient of determination, is 0.0903.